OpenStudy (anonymous):

Please Help!!! D= r*t??? For his morning workout, a boxer bicycles 8 miles and then jogs back to his house along the same route. If he bicycles 6 miles per hour faster than he jogs and the entire workout was 2 hours, how fast does his jog? a) Declare Variable b) Equation c) Solving d) Answer -------------------------------------------- So far I have a) x: For how fast the boxer jogs x+6: For how fast the boxer bicycles Now this is where I get fuzzy is it Distance= Rate * Time Plane 16 x+6 2 Car 16 x 2

zepdrix (zepdrix):

Hey Binky c: Hmm what is this Plane/Car stuff about?

OpenStudy (anonymous):

OH! This reminds me of something actually thank you! I knew it looked funny

zepdrix (zepdrix):

Oh I didn't plug in the distance, hehe..

zepdrix (zepdrix):

bahh i dunno >.< my brain so fried today..

OpenStudy (anonymous):

who would jog for freaken 2 hrs?? it's suicidal!!!! D:

OpenStudy (anonymous):

Well I think so So the variables it would be x: How fast the boxer jogs Distance = Rate * Time Bike 8 x+6 8/(x+6) Jog 8 x 8/x So the equation would be 8/x=8/(x+6) + 2

OpenStudy (anonymous):

@zepdrix haha so is mine now... I thought this problem was all finished but my friend that helped me with it earlier this weekend texted me and said she had some doubts and maybe I should look at it again. SO that lead to a whole chain of problems and uncertainty.

OpenStudy (anonymous):

@sourwing No kidding haha I am good for an hour, but two hours Im pretty dead.

OpenStudy (anonymous):

@BinkyBoo , it should be 8/x + 8/(x+6) = 2 x = 6 mi/hr

OpenStudy (anonymous):

so \[\frac{ 8 }{ x } + \frac{ 8 }{ x+6 }=2\] \[x(x+6) [\frac{ 8 }{ x}+\frac{ 8 }{ x+6 }]= [2] x(x+6)\]\[x(x+6)(\frac{ 8 }{ x }) + x(x+6)(\frac{ 8 }{ x+6})= 2 (x)(x+6)\]\[8(x+6) +8x=2x(x+6)\] \[8x+48+8x=2x ^{2}+12x\]\[16x+48=2x ^{2}+12x\] \[0=2x ^{2}-4x-48\]

OpenStudy (anonymous):

\[0=2(x ^{2}-2x-24)\] \[0=2(x+4)(x-6)\] 0=2 or x+4=0 or x-6=0 can't be x=-4 x=6 Cant be can be

OpenStudy (anonymous):

so x= How fast the boxer jogs is 6 mph ???

OpenStudy (triciaal):

yes

OpenStudy (anonymous):

Thank you!